7(1 divided by 7)^7-t=7
Answer by checkley77(12569)

7(1/7)^7-t=7
7*1/823,543-T=7
7/823,543-T=7
-T=7-7/823,543
-T=7-.0000085
-T=6.999915
T=-6.999915 ANS.